OpenStudy (anonymous):

Show your work to prove that the equation v = -4t - 2 correctly represents the function for the data in this table A t v table shows the following coordinate pairs: negative two comma six, negative one comma two, and zero comma negative two.@HELP!!!! @tkhunny

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (tkhunny):

(-2,6) (-1,2) (0,-2) (t,v) Have you considered substituting the values for 't' and seeing if you manage the corresponding 'v'?

OpenStudy (anonymous):

Show your work to prove that the equation v = -4t - 2 correctly represents the function for the data in this table.

OpenStudy (tkhunny):

Have you considered substituting the values for 't' and seeing if you manage the corresponding 'v'? Let's try t = 0 and see if we get v = -2. -4(0) - 2 = 0 - 2 = -2. which is, indeed, -2.